# Low Regularity of Self-Similar Solutions of Two-Dimensional Riemann problems with Shocks for the Isentropic Euler system

### v1

@@ Appendix A DiPerna-Lions-Type Commutator Estimates
Then we complete the proof of (A.3) for $I_{\varepsilon,4}$ similar as for $I_{\varepsilon,1}$ above, by using the $L^{p}$–estimate the difference quotient for the $b$-term, and the $L^{q}$–estimate of the shift for the $u$-term. Then estimate (A.2) holds.
@@
Acknowledgements. The research of Gui-Qiang G. Chen was also supported in part by the UK Engineering and Physical Sciences Research Council Award EP/L015811/1, EP/V008854, and EP/V051121/1. The research of Mikhail Feldman was supported in part by the National Science Foundation under Grants DMS-2054689 and DMS-2219391, and the Steenbock Professorship Award from the University of Wisconsin-Madison. The research of Wei Xiang was supported in part by the Research Grants Council of the HKSAR, China (Project No. CityU 11304820, CityU 11300021, CityU 11311722, and CityU 11305523), and in part by the Research Center for Nonlinear Analysis of the Hong Kong Polytechnic University. For the purpose of open access, the authors have applied a CC BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.

### v2

@@ Appendix A DiPerna-Lions-Type Commutator Estimates
Then we complete the proof of (A.3) for $I_{\varepsilon,4}$ similar as for $I_{\varepsilon,1}$ above, by using the $L^{p}$–estimate of the difference quotient for the $b$-term, and the $L^{q}$–estimate of the shift for the $u$-term. Then estimate (A.2) holds.
@@
Acknowledgements. The research of Gui-Qiang G. Chen was supported in part by the UK Engineering and Physical Sciences Research Council Awards EP/L015811/1, EP/V008854, and EP/V051121/1. The research of Mikhail Feldman was supported in part by the National Science Foundation under Grants DMS-2054689 and DMS-2219391, and the Steenbock Professorship Award from the University of Wisconsin-Madison. The research of Wei Xiang was supported in part by the Research Grants Council of the HKSAR, China (Project No. CityU 11304820, CityU 11300021, CityU 11311722, and CityU 11305523), and in part by the Hong Kong Polytechnic University internal grant (No. P0045335). For the purpose of open access, the authors have applied a CC BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
@@
Funding Declaration: The research was supported in part by the UK Engineering and Physical Sciences Research Council Awards EP/L015811/1, EP/V008854, and EP/V051121/1; the US National Science Foundation under Grants DMS-2054689 and DMS-2219391, and the Steenbock Professorship Award from the University of Wisconsin-Madison; and the Research Grants Council of the HKSAR (China) Project Nos. CityU 11304820, CityU 11300021, CityU 11311722, and CityU 11305523, and the Hong Kong Polytechnic University internal grant (No. P0045335).